OpenStudy (anonymous):

Using your function f(x) from question 2, demonstrate and explain how to find the average rate of change between year 3 and 5, and between year 5 and 7. Explain what the average rate of change represents to the frog population.

OpenStudy (anonymous):

my function is f(x) = a times b^x

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@wmj259

OpenStudy (wmj259):

what was the question 2 function?

OpenStudy (anonymous):

f(x) = a * (b)^x

OpenStudy (anonymous):

are you there? @wmj259

OpenStudy (calculusfunctions):

The average rate of change (slope of a secant) of a function f(x) over the interval\[x _{1} \le x \le x _{2}\]is\[\frac{ f(x _{2})-f(x _{1}) }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

thank you!!! but what is x2 and x1????

OpenStudy (anonymous):

@calculusfunctions ^^

OpenStudy (calculusfunctions):

x sub 1 is the smaller of the two x values and x sub 2 is the larger.

OpenStudy (calculusfunctions):

Hence if you have to find the average rate between year 3 and year 5, then x sub 1 is 3 and x sub 2 is 5. Understand @haha.lena ?

OpenStudy (anonymous):

yes! thank you so much! you are officially medalled. lol xD

OpenStudy (calculusfunctions):

lol You're welcome and thank you!